OpenStudy (anonymous):

A man bought 42 stamps, some 13¢ and some 18¢. How many of each kind did he buy if the cost was $6.66? If x represents the number of 13 cent stamps and y the 18 cent stamps, which system represents the problem?

OpenStudy (anonymous):

let x+y=42, then y = 42-x so according to question, 0.13x + 0.18(42-x) = 6.66